OpenStudy (anonymous):

Which of the following is the reciprocal of x2 – 2x – 8?

OpenStudy (anonymous):

A. (x + 2)(x – 4) B. 1 (x + 2)(x – 4) C. 1 (x – 8) D. 1 (x + 2)

OpenStudy (anonymous):

Well, What does a reciprocal mean?

OpenStudy (anonymous):

What would the reciprocal of say 10 be?

OpenStudy (anonymous):

0.1

OpenStudy (anonymous):

Right. So when you foil \(\ \huge \color{blue} x^2-2x-8 \), what do you get?

OpenStudy (anonymous):

Reciprocal is 1/x. \[\frac{1}{x^2-2x-8}\] Factoring that out; \[\frac{1}{(x-4)(x+2)}\] Therefore B

OpenStudy (anonymous):

\(\ \huge (x+2)(x-4) \). You can also write that as a reciprical: \(\ \huge \frac{1}{(x+2)(x-4)} \). So, the correct answer is B. Do you understand that?